### v2

# The Surprising Effectiveness of Multimodal Large Language Models for Video Moment Retrieval

Abstract
Recent studies have shown promising results in utilizing multimodal large language models (MLLMs) for computer vision tasks such as object detection and semantic segmentation. However, many challenging video tasks remain under-explored. Video-language tasks necessitate spatial and temporal comprehension and require significant compute. Therefore, prior works have developed complex, highly specialized architectures or leveraged additional input signals such as video transcripts to best encode contextual and temporal information, which limits their generality and can be impractical. One particularly challenging task is video moment retrieval, which requires precise temporal and contextual grounding. This work demonstrates the surprising effectiveness of leveraging image-text pretrained MLLMs for moment retrieval. We introduce Mr. BLIP[Uncaptioned image] (Mr. as in Moment Retrieval), a multimodal, single-stage model that requires no expensive video-language pretraining, no additional input signal (e.g., no transcript or audio) and has a simpler and more versatile design than prior state-of-the-art methods. We achieve a new state-of-the-art in moment retrieval on the widely used benchmarks Charades-STA, QVHighlights, and ActivityNet Captions. Notably, we attain over 9% (absolute) higher Recall (at 0.5 and 0.7 IoU) on the challenging long-video multi-moment QVHighlights benchmark. Our code is publicly available  [^1] https://github.com/sudo-Boris/mr-Blip.
1 Introduction
Figure 1: Advantages of Mr. BLIPRefer to caption over prior work. In this work (top), we cast the task of predicting relevant windows based on a natural-language query as a sequence-to-sequence task. We utilize the frames directly and leverage an MLLM as a general-purpose interface [21] to predict a variable number of relevant moments that represent the query. In contrast, prior works (bottom) train static, task-specific multi-proposal prediction heads on top of a feature extractor to predict a fixed set of candidates and then sort the predicted moments based on a confidence score.
The recent success of pretrained large language models (LLMs) [3, 65, 14] has inspired the development of generative image-text pretrained multimodal large language models (MLLMs) [1, 36, 23] that can comprehend vision and language modalities jointly. Furthermore, recent studies have demonstrated that large language models can be used as general-purpose interfaces [21] and can be adapted for image-based computer vision tasks such as object detection [9], image captioning [60], or visual grounding [47, 66]. However, due to higher computational and annotation costs, large-scale pretraining on video data is more challenging to scale. To circumvent the issue, recent studies leverage image-text pretrained models for image-to-video transfer learning [61, 62, 35, 42, 16, 25]. While such models offer promising results in the direction of video-text retrieval, video captioning, or multiple choice video question answering, generative MLLMs have not been extensively explored for video-based moment retrieval.
Video moment retrieval requires the precise temporal localization of all moments associated with an open-ended natural-language query in an untrimmed video. This task has multiple potential valuable applications, such as video search or indexing. More concretely, it requires understanding, discrimination, and temporal localization of multiple events in potentially minutes-long videos given an open-ended natural-language query. In the context of MLLMs, it is an open question of how to best model this task as a sequence-to-sequence prediction task, e.g., how to best represent timestamps so the model can reason about them and predict start and end timestamps correctly.
As illustrated in Fig. 1 (bottom), recent works generally approach the challenge of moment retrieval by leveraging an image-text aligned feature fusion module and train additional task-specific prediction heads that propose a fixed set of potentially relevant moment candidates. The final set of relevant moments is then selected from all candidates based on their confidence scores by picking the top-$k$ windows given $k$ target windows in a benchmark. In practice, this approach is unsatisfactory, as selecting the right number of relevant windows given a fixed set of candidates is not trivial. Moreover, these methods often require expensive video pretraining to interpret the video modality.
In contrast, to the best of our knowledge, we are the first to cast the moment retrieval task as an open-ended sequence-to-sequence problem, leveraging a generative MLLM and demonstrating the surprising effectiveness of its contextual understanding abilities. This approach has two significant advantages. (1) We avoid the constraint of generating a fixed set of moment proposals (Fig. 1, bottom). Our model can predict a variable number of intervals that reflect the actual number of relevant moments, eliminating the ambiguity of selecting a subset of relevant moments given a fixed number of candidates (Fig. 1 top). (2) By leveraging an image-text pretrained MLLM we do not require expensive video pretraining and can quickly adapt to understanding the video modality.
In this work, we develop Mr. BLIP[Uncaptioned image], an image-text pretrained, multimodal large language model based on BLIP-2 [36] and a novel multi-modal input sequence. The model design is capable of interpreting video data, and leverages interleaved temporal grounding of the input video signal (in the form of time-stamps) to enhance its predictions. As illustrated in Fig. 1, we also start of with both visual frames and query. Prior work (bottom), often leverages video features [19] and CLIP [49] to then train a complex, task-specific feature fusion module that predicts a fixed set of candidates. In contrast, Mr. BLIP (top) leverages a sampled subset of the frames directly and utilizes the MLLM’s attention mechanism to learn the contextual and temporal relationship between visual frames, their interleaved temporal grounding, and the respective natural language query. Later in the paper, we discuss the relevance of deliberate design choices by conducting extensive ablation studies and demonstrate the surprising effectiveness of leveraging image-text pretrained MLLM for moment retrieval. Our approach achieves state-of-the-art results across the widely used benchmarks, Charades-STA [20], QVHighlights [33], and ActivityNet Captions [29].
In summary, our main contributions are the following: (i) We introduce Mr. BLIP for moment retrieval, leveraging image-text pretrained MLLMs by casting the moment retrieval task as an open-ended sequence-to-sequence problem. (ii) To enhance the temporal understanding of events in input videos, we design a novel multimodal input sequence. (iii) Our Mr. BLIP model improves the state-of-the-art across the most widely used moment retrieval benchmarks [20, 33, 29]. (iv) Finally, extensive experiments and ablations demonstrate the effectiveness of Mr. BLIP and its design choices. We find that representing time as absolute seconds and interleaving it with the respective frames yields the best performance.
2 Related Work
2.1 Moment Retrieval Models
Moment Retrieval analyzes an untrimmed video and aims to find the relevant clip for a given open-ended natural language query. Conventional approaches fall into either proposal-based or proposal-free methods. Proposal-based methods learn to identify moment candidates given predefined proposals, e.g., sliding windows [2, 20] and temporal anchors [8, 53], in a first stage. In a second stage, these candidates are further refined to better match the text query.
Proposal-free works include regression-based methods [33, 58, 39, 46, 63] that involve learning the interactions between both modalities to directly predict the temporal boundaries of a relevant moment without the need for a first stage to generate initial proposals. Nevertheless, proposal-free methods still predict a fixed number of candidate windows alongside a confidence score.
The predicted windows are sorted based on the confidence score and are evaluated with proxy metrics based on different overlapping thresholds such as Recall@{0.5, 0.7}.
With the adoption of transformers and the success of the query-based design for object detection (DETR), recent works also employ this design for the moment retrieval task. Due to the slow and inefficient learning ability of moment queries, recent works have focused on initializing the queries with additional prior information [24] or developing complex attention modules [44, 45]. One of the state-of-the-art models, EaTR [24], utilizes pseudo event windows, obtained from the temporal self-similarity matrix (TSM) [26] and feeds them as positional queries. Conversely, our proposed design predicts moments without relying on any prior information. Furthermore, these models [44, 45] leverage 3D-temporal features such as Slow-Fast [19] or C3D [51] to predict a fixed set of moment candidates, as illustrated in Fig. 1 bottom. In contrast, we extract the individual frame features directly, omitting the reliance on expensive video-level features.
On the other hand, Yan et al. [58] leverage the large-scale image-text pretrained CLIP [49] model for a unified video localization model. CLIP is used to encode video frames and text queries that are then concatenated and fed into a novel video-text fusion module based on the transformer [52] architecture with multiple prediction heads. This approach appears promising because it leverages language priors from the pretrained CLIP model early on in their model. Nevertheless, their approach requires expensive video pretraining on the Kinetics400 [27] and Kinetics700 [6] datasets. Finally, each prediction head is defined to solve a separate task: moment retrieval, action segmentation, and temporal action localization, respectively. In contrast, we leverage the strong contextual understanding ability of MLLMs by casting the moment retrieval task as an open-ended language modeling task and do not require any video pretraining. Similar to Yan et al. [58], our design also combines the language and vision context early on in the model so that the generation process is conditioned on both modalities from the beginning. Moreover, Luo et al. [41] also utilize image-text pretrained models but differ in methodology to Mr. BLIP. Their approach is metric learning based, focusing on moment retrieval by matching video and text, whereas ours is regression-based, directly predicting moment boundaries.
2.2 Computer Vision Tasks as Language Modeling
Recent works [47, 59, 12, 10, 9, 11, 13, 28, 37, 55, 60, 66] have demonstrated the possibility of casting various computer vision problems as language modeling tasks, addressing object detection [9, 48], semantic segmentation [30], or other vision-centric tasks [56]. In this work, we also cast temporal localization as a language modeling task. Unlike most previous work that focuses on image-level localization, similar to Yang et al. [59], we address the challenge of event localization in untrimmed videos. However, we focus on moment retrieval, i.e., detecting moments in a video given a query, and not on dense event captioning.
Yu et al. [62] leverage BLIP-2 [36] for a naive approach to moment retrieval by iterating over the sampled subset of video frames and prompting the multimodal large language model to classify each frame as relevant/irrelevant to the query of their downstream task. Yu et al. do not achieve state-of-the-art performance on the QVHighlights [33] but they hint at the possibility of leveraging MLLMs for moment retrieval tasks. Here, we explore this idea and extend it. We incorporate multiple frames and interleave their temporal grounding when constructing the input sequence to adapt BLIP-2 to understand their contextual and temporal relationship. We further fine-tune our model to output the retrieved moments in an open-ended sequence-to-sequence fashion and not only evaluate the probabilities of two tokens for binary classification (as in Yu et al.).
2.3 Image-to-Video Transfer Learning
The limited public availability of high-quality large-scale video-language datasets and immense compute requirements pose a significant challenge for large-scale video-language pretraining. The idea of leveraging image-language pretrained models for image-to-video transfer learning by utilizing a limited number of video frames to enhance learning efficiency has proven effective as evidenced by many recent publications [61, 62, 5, 16, 17, 25, 35, 42, 43, 34, 57, 54].
Yu et al. [61] leverage BLIP-2 [36] and the in-context learning abilities of MLLMs to fine-tune a model on egocentric videos using a fixed number of frames. An image-language pretrained MLLM allows them to train their model without requiring massive, naturalistic egocentric video datasets and transfer the image-language knowledge to the video-language modality.
Yu et al. [62] also make use of the contextual understanding of MLLMs to perform video question-answering tasks. The authors train a model to individually select 4 question-relevant frames and then prompt BLIP-2 using the separately obtained frame embeddings as context. The authors argue that performance decreases when providing more frames in the case of video question answering. We also leverage the pretrained BLIP-2 model and fine-tune it on a downstream video-language task, moment retrieval, but demonstrate strong performance improvements when providing significantly more frames as context.
3 Mr. BLIP[Uncaptioned image] for Moment Retrieval
Figure 2: Mr. BLIPRefer to caption model overview. We leverage a pretrained MLLM based on an encoder-decoder LLM such as BLIP-2 [36]. We sample F frames (left) and provide them separately as input to our LLM encoder $\mathcal{E}$ (right). We construct the prompt for the LLM by concatenating the frame embeddings, the timestamps of each sampled frame, the video duration, the moment retrieval query, and a task prompt. (We do not visualize the task prompt for illustration purposes.) We finetune parts of the LLM leveraging parameter-efficient finetuning [22]. The LLM decoder $\mathcal{D}$ outputs a sequence of potentially multiple retrieved moments by predicting the global BOS and EOS tokens, denoted by ##\##, the start of window and end of window tokens denoted as #$\#_{start}$ and #$\#_{end}$, and the respective start and end times for each window.
The task of moment retrieval is to temporally localize all relevant moments in an untrimmed video given an open-ended natural language query. Therefore, a key challenge is to effectively model the contextual and temporal relationship between the different events in the video, for example, as illustrated in Fig. 2, a man is talking to the camera followed by him watching a show. Furthermore, due to the density of information in a video, it is computationally unfeasible to leverage all frames of a video as context in a single pass-through model.
To tackle these challenges, we cast moment retrieval as a language modeling task and leverage the contextual understanding ability of generative MLLMs to interpret and comprehend the semantic and temporal action development in frames that represent a video. We first design an MLLM-based multimodal model, Mr. BLIP, that can interpret video frames and their respective temporal grounding in the video and learn to associate a moment of the video to a natural language query (Section 3.1 and Fig. 2). Second, we develop a training strategy that effectively utilizes the limited amount of training data (Section 3.2).
3.1 Model
Our work explores the possibility of leveraging MLLMs capable of interpreting visual data for moment retrieval in videos given a natural language query. We design a model that can capture contextual and temporal relationships between the actions in an untrimmed video and the query. To do so, we cast the traditional moment retrieval task as an open-ended sequence-to-sequence problem, where we design a novel multimodal input sequence that contains (a) the visual semantic context given the sampled frames of the video, (b) the temporal context for each action given by the temporal grounding of each frame, and (c) the query in the form of a natural language description. The output of our model is a time interval sequence where each time interval represents a moment associated with the query and follows the formatting of a nested list containing potentially multiple moments with a start and end time.
As illustrated in Fig. 2, our architecture takes as visual input $F$ sub-sampled video frames. There are multiple ways of modeling the temporal context for both input and output, which significantly affects the downstream performance. We find that modeling the timestamps as absolute seconds and, as visualized, interleaving them with the frames yields the best performance. With this representation, it is likely easier for the model to associate each frame $f_{i}$ with its respective timestamp $t_{i}$.
3.1.1 BLIP-2 Architecture.
We leverage the contextual understanding capabilities of MLLMs to interpret video data and detect relevant moments given an open-ended natural language query. To best leverage both the visual and the language signal, we adopt BLIP-2 [36] as the multimodal backbone for our model. BLIP-2 is a recent state-of-the-art MLLM pre-trained on a large image-text corpus. We leverage BLIP-2 to bootstrap the learning on video data, given its limited availability and expensive compute requirements for training. BLIP-2 comprises (1) a frozen image encoder [15, 18] combined with (2) a Q-Former, which is a trainable transformer [52] module that functions as a translation unit between the image encoder and the last component, and (3) a frozen LLM [14, 65].
3.1.2 Multimodal sequence construction.
To model contextual and temporal relationships between the different actions in a video and a natural language query, we cast moment retrieval as a sequence-to-sequence task and develop a novel multimodal input sequence. It consists of frames $f_{n}$, timestamps $t_{n}$ ($n=1,...,F$), video duration $d$, query $q$ and task prompt $p$ stating the task at hand as input. The task prompt we use is “Given the video and the query, find the relevant windows. Relevant windows: ”. Finally, the model is tasked to predict the sequence of potentially multiple relevant windows $m$, which follows the formatting of a nested list of moments with a start and end time in seconds: $y=[[t_{start}^{1},t_{end}^{1}],[t_{start}^{2},t_{end}^{2}],...]$.
Notably, there are multiple ways of representing timestamps $t_{n}$ and thus designing the multimodal sequence. Time can be represented in absolute or relative form, e.g., “79.9” seconds or “0.40” for a relative position; it could be a decimal number or be rounded to the nearest full integer value, e.g., “80” seconds or “40” for a relative position; the timestamps could be concatenated after the frames $f_{n}$ or interleaved with them. As illustrated in Fig. 2, we find that representing time as simple text tokens of whole seconds (integer absolute form) and interleaving them with corresponding frames yields the best performance. We hypothesize this is the case because of how decimal numbers are tokenized. A decimal number, such as “79.9”, is tokenized into three tokens, whereas integers (up to “100”) are tokenized as a single token. One timestamp token corresponds to one frame. Additionally, placing each timestamp next to its associated frame aids the model to learn better which time token corresponds to which frame. In Sec. 4.2, we explore the different variations of representing time and concatenation of the different sequence elements of the input.
We find that representing timestamps as whole seconds (rounding decimal numbers to the nearest integer) and interleaving them with the frames yields the best performance. This finding leads to our final design for our multimodal input sequence $x=[f_{1},round(t_{1}),f_{2},round(t_{2}),...,f_{F},round(t_{F}),d,q,p]$.
3.1.3 Frame Encoder.
We denote BLIP-2’s frozen image encoder combined with Q-Former as our general-purpose frame encoder. The frame encoder is separately applied to each of the $F$ sub-sampled frames, generating the frame embeddings and projecting them into the language space, yielding a shared embedding space. Prior works [59, 64] incorporate a further trainable transformer-based module to generate temporally and contextually correlated frame- or video-level embeddings. In contrast, we solely leverage the self- and cross-attention mechanisms of our LLM to learn the temporal and contextual relationships between frames.
3.1.4 Text Decoder.
For the text decoder, we have chosen an encoder-decoder-based LLM. We leverage the instruction-trained Flan-T5 model [14] since the pretrained BLIP-2 frame encoder is aligned to it. The LLM encoder $\mathcal{E}$ learns the correlation between the input sequence elements $x=[f,t,d,q,p]$ due to its bi-directional attention mechanism and generates new multimodal contextualized output embeddings. The text decoder $\mathcal{D}$ is based on a causal transformer decoder that cross-attends to the encoder outputs $x$. At each autoregressive step $i$, the decoder self-attends to the previously generated tokens $\hat{m}_{<i}$ to output a conditioned probability distribution over the vocabulary of text: $m_{i}=\mathcal{D}(\hat{m}_{<i},\mathcal{E}(x))\in\mathbb{R}^{V}$, where $V$ is the size of the vocabulary.
3.2 Training
In this section, we describe how we train our generative MLLM on the task of moment retrieval that has been cast from a traditional computer vision problem to an open-ended natural language sequence-to-sequence task.
3.2.1 Parameter-Efficient Finetuning.
To train our model, we optimize for the standard maximum likelihood objective. Given the task input $x$ and the decoder target text $y$, the objective is to minimize the following loss:
$\mathcal{L}(x,y)=-\sum^{L-1}_{i=1}\;log\;p_{\theta}(y_{i+1}\,|\,x,y_{1:i}),$ (1)
where $L$ is the length of the decoder target sequence and $p_{\theta}$ denotes the output probability distribution over the vocabulary of text.
To further optimize training and best leverage the limited video data available, we implement parameter-efficient finetuning techniques, specifically LoRA [22], and randomly sample video frames. This allows us to only train roughly 20M parameters of the LLM, which is significantly less than in prior work [59, 58]. Additionally, by randomly sampling the frames, we optimize those weights by analyzing different frames and timestamps for the same videos each epoch. For further details on the random sampling, please refer to Appendix 0.A.1.
3.3 Inference
At inference, we uniformly sample video frames to have the best temporal distribution. The text decoder then autoregressively generates the moment sequence by sampling from the model likelihood. In practice, we use beam search following prior work [59]. Finally, the natural language moment sequence is converted into a nested list of predicted moments interpreting the sequence as a common Python list, automatically fixing a few common formatting mistakes.
4 Experiments
This section describes the effectiveness of our design choices and compares our method to the state-of-the-art. In Section 4.1, we first elaborate on our experimental setup and implementation details. We then discuss our ablation studies in Section 4.2 and present our comparison to the state-of-the-art in Section 4.3. Finally, in Section 4.4, we show qualitative results from our approach.
4.1 Experimental Setup
4.1.1 Benchmarks.
We validate Mr. BLIP on the three most widely used video moment retrieval (MR) datasets:
Charades-STA [20] includes 9,848 videos with an average duration of 30.6 seconds. The dataset contains 16,128 annotations with an average moment length of 8.1 seconds and an average query length of 7.22 words. The dataset is originally divided into two splits: training (12,408) and test (3,720). To avoid overfitting on the test set during training, we designate a part of the training set as a new validation set. The videos used in the validation set are not contained in the training set. Our new dataset split consists of a split for training, validation, and testing, with 11,166, 1,242, and 3,720 annotations, respectively. For a fair comparison, after completing our ablations, we train our final model on the original training set and report numbers on the test set.
QVHighlights [33] is one of the most recent benchmarks and contains 10,148 videos with a duration of 150 seconds. The videos were cropped out of YouTube videos, and each video is annotated with at least one query with an average length of 11.3 words describing the relevant moment. The target windows have an average length of 24.6 seconds. The dataset is split into training, validation, and test sets with 7,218, 1,150, and 1,542 queries, respectively. This benchmark is challenging because one query can be associated with multiple moments in a video. The test set targets are withheld and thus guarantee a fair benchmark. The evaluation for the test split can only be measured through submitting the prediction to the evaluation server [^2] Evaluation server, QVHighlights: https://codalab.lisn.upsaclay.fr/competitions/6937.
ActivityNet Captions [29] contains 20,000 videos with an average duration of 2 minutes. The dataset contains 72,000 segments where each was human-annotated with a caption that contains on average 13.5 words. The dataset is divided into three splits, train (37,421), val_1 (17,505), and val_2 (17,031). Following [58], we use the train split for training, val_1 for validation, and val_2 for testing.
4.1.2 Metrics.
The most commonly used metrics for MR are Recall@K and mean average precision (mAP) computed under different Intersection over Union (IoU) thresholds. The Recall@K metric is defined as the percentage of the top-K predicted segments having a larger temporal IoU than the threshold with a ground truth segment. Following the recent development [32, 33, 44, 45], we report the more challenging $R1@0.5$ and $R1@0.7$ scores, which correspond to the $Recall@1$ scores at IoU thresholds of $0.5$ and $0.7$, respectively, and the mean IoU (mIoU) score. Following prior work, for MR on QVHighlights, we report the mAP score at IoU thresholds of $0.5$ and $0.75$, respectively.
4.1.3 Implementation details.
Considering the relatively small scale of the datasets, we choose not to finetune the entire 3 billion parameter-LLM backbone of our model but rather leverage parameter-efficient finetuning [22] to train only 19 million parameters, as described in Section 3.2.
Since our objective is a generative language modeling task, our model can output minor formatting inconsistencies that would invalidate the prediction if not treated otherwise. Similar to [59], we post-process the model’s outputs by applying heuristics to clean up minor flaws in the prediction.
For Charades-STA, we extract 20 video frames for each video, which is significantly less than prior work which requires 128 frames [58]. We train for up to 50 epochs with a batch size of 32 samples split across 4 A100-80GB GPUs for about 55 GPU hours. For QVHighlights and ActivityNet Captions, we extract 60 video frames for each video, which is less than required by prior work. We train for up to 50 epochs with an effective batch size of 32 samples by splitting them across 8 A100-80-GB GPUs and leveraging accumulated gradients set to 4 (i.e., we have one sample per GPU), and train for about 170 GPU hours.
We start with a learning rate (lr) of 1e-8, perform linear lr warmup to 3e-4 for 10% of the total number of iterations and then apply a cosine lr decay. We utilize the AdamW [40] optimizer and sample the frames randomly during training.
4.2 Ablation Studies
Table 1: Ablation on design of prompt. Evaluation on Charades-STA.
Input components R1@0.5 R1@0.7 mIoU
Frames only 43.89 23.47 41.25
+ Duration 58.28 34.00 51.01
+ Frame timestamps 67.28 45.26 57.24
+ Task description 67.28 46.7 57.46
By default, our Mr. BLIP models for Charades-STA (QVHighlights/ ANet Captions) leverage 20 (60) visual frames, their associated timestamps, video duration, a query, and a task prompt. If not stated otherwise, we represent time as simple text tokens of whole seconds and interleave these with frame tokens. In the following, we ablate the impact of these design choices on the downstream moment retrieval performance by reporting results on our Charades-STA validation set.
4.2.1 Design of sequence.
In Table 1, we analyze the effectiveness of each part of the multimodal input sequence of our MLLM, by cumulatively adding each component starting with only providing the frames. The query is always part of the sequence. Adding the video duration, in addition to the visual inputs, improves the downstream performance significantly (row 2 vs. row 1). This shows the importance of providing the model a point of reference to infer the temporal location of each frame and demonstrates the ability of the MLLM to learn this association. Furthermore, additionally providing the specific timestamps at which each frame was sampled improves the model performance even further by a significant margin (row 3 vs. row 2). This demonstrates the relevance of designing the input context to provide as much relevant information as possible, while still not relying on additional computation such as generating and leveraging transcripts. Finally, we surprisingly find that adding a task prompt describing what we expect the model to do ("Given the video and the query, find the relevant windows. Relevant windows: ") improves model performance as well (row 4 vs. row 3) and might be worth exploring in future work. We hypothesize this is due to the instruction tuning in our leveraged Flan-T5 [14] model.
Table 2: Design of timestamps. We demonstrate the importance of how we represent time: (A) Representing time as decimal numbers has the lowest performance while providing absolute timestamps improves the performance over relative temporal positions. (B) Interleaving frames with their associated timestamps yields further significant improvement. The ablation was done on the Charades-STA validation set.
# Representation Precision Interleaved R1@0.5 R1@0.7 mIoU
(1) Relative decimal 62.30 37.22 54.03
(2) Absolute decimal 62.38 36.33 53.59
(3) Relative integer 63.10 36.01 53.86
(4) Absolute integer 64.39 41.80 56.00
(5) Relative integer 65.19 44.94 56.99
(6) Absolute integer 67.28 46.70 57.46
4.2.2 Design of timestamps.
A central question for this work is how MLLMs can best represent and reason about timestamps, as they have to be interpreted in the input and predicted in the output. We compare different options for representing the timestamps and their impact on the final downstream model performance. We explore using relative positions w.r.t the video duration vs. representing each timestamp as absolute time. In both cases, we explore representing the numbers as decimal numbers, e.g., “79.9” seconds or “0.40” for a relative position, and whole integers, e.g., “80” seconds or “40” for a relative position. In these experiments, the format of the timestamps in model input and output is always consistent. In the case of relative positions, we post-process the output to yield an absolute value to compare to the ground truth for the final evaluation.
In Table 2 we observe, that that representing the timestamps in decimal form results in lower performance for both absolute and relative form, see rows (1,2) vs. rows (3, 4). We hypothesize this is the case because of how decimal numbers are tokenized. A decimal number, such as “79.9”, is tokenized into three tokens, whereas integers (up to “100”) are all tokenized as one token. This shows the importance of designing the input and outputs with the representation of the MLLM in mind. Further, we observe that absolute position in seconds yields better performance than the relative representation when using integer precision.
Finally, we explore the design choice of ordering frame and time tokens. In rows (1) through (4), we concatenate all frames together, followed by a concatenation of the timestamps, which are each separated by a separator token (“>”). In rows (5) and (6), we demonstrate the benefit of interleaving frame tokens with time tokens so that each frame is followed by its respective timestamp representation (as illustrated in Fig. 2). This again shows the importance of a careful prompt design. We hypothesize that the attention mechanism can more easily learn the correlation between a frame and its associated timestamp.
Table 3: Ablation on the number of input frames. Evaluation on Charades-STA eval.
#Frames R1@0.5 R1@0.7 mIoU
5 54.50 32.72 48.26
10 64.63 41.80 55.81
15 67.79 44.07 57.71
20 67.28 46.70 57.46
25 67.95 43.35 57.91
30 64.58 41.51 55.79
4.2.3 Number of input frames.
In Table 3, we show the impact of the number of input frames and demonstrate the ability of our image-text pretrained MLLM to adapt to the video modality.
For Charades-STA with an average video duration of 30 seconds, using 20 frames achieves the highest performance on R1@0.7, improving over 15 frames by $2.63\%$ and over 25 frames by $2.25\%$, respectively. We observe, that inputting 25 frames does improve the R1@0.5 and mIoU scores compared to leveraging 20 frames, but only by $0.67\%$ and $0.45\%$, respectively while increasing compute requirements. Leveraging 30 frames reduces the performance, albeit theoretically adding more information.
The number of input frames is a hyperparameter that we explore for each dataset on the validation set. For QVHighlights, which consists of 150-second long videos, and for ActivityNet, which on average has 120-second long videos, we have found 60 frames to perform the best, showing the ability of Mr. BLIP to comprehend a relatively long sequence of interleaved visual and textual tokens.
4.3 Comparison to the State-of-the-Art
Table 4: Comparison with state-of-the-art. Charades-STA [20], QVHighlights [33], and ActivityNet Captions [29]. CLIP: UnLoc [58] pretrains the backbone with action classification datasets Kinetics 400/700. SF: Slow-Fast backbone. Discussion in the main text.
Method Backbone mIoU R1@.5 R1@.7 mAP@.5 mAP@.75
QVHighlights [33] Validation set
EaTR [24] I3D 61.36 45.79 61.86 41.91
UnLoc-L [58] CLIP 66.10 46.70
Mr. BLIP[Uncaptioned image] BLIP-2 76.13 63.35 69.39 55.78
QVHighlights [33] Test set
SeViLa [62] BLIP-2 54.50 36.50
UniVTG [38] SF+CLIP 58.86 40.86 57.60 35.59
QD-DETR [45] SF+CLIP 62.40 44.98 62.52 39.88
CG-DETR [44] SF+CLIP 65.43 48.38 64.51 42.77
Mr. BLIP[Uncaptioned image] BLIP-2 74.77 60.51 68.12 53.38
Charades-STA [20] Test set
UniVTG [38] SF+CLIP 50.10 58.01 35.65
QD-DETR [45] SF+CLIP 57.31 32.55
CG-DETR [44] SF+CLIP 50.13 58.44 36.34
UnLoc-L [58] CLIP 60.80 38.40
EaTR [24] I3D 68.47 44.92
Mr. BLIP[Uncaptioned image] BLIP-2 58.63 69.31 49.29
ActivityNet Captions [29] Single sentence MR. Test set
VLG [50] GCNeXt 46.30 29.80
UnLoc-L [58] CLIP 48.30 30.20
Mr. BLIP[Uncaptioned image] BLIP-2 53.92 35.55
Next, we compare Mr. BLIP to the state-of-the-art (SoTA) approaches for MR following the best design choices derived from our ablations in Section 4.2.
4.3.1 Moment Retrieval
As outlined in Table 4, for QVHighlights, Mr. BLIP significantly outperforms the previous SoTA [44] on all metrics. Our model improves the performance by 9.21%, 11.41%, 3.9%, and 10.18% for R1@0.5, R1@0.7, mAP@0.5, and mAP@0.75, respectively. Interestingly, we find that the improvement on the more difficult metrics (with a higher IoU threshold) is higher than on the easier metrics (with a lower IoU threshold). This implies that our model may not only predict significantly more acceptable windows, i.e., windows with an IoU>0.5, but a larger subset of those have a higher precision, indicated by the even larger improvements for IoU>0.7 and IoU>0.75, respectively.
Notably, we also outperform the best reported approach on QVHighlights validation set, UnLoc-L[58], with far fewer input frames, 128 vs. 20, and without expensive video pretraining of the backbone.
As shown in Table 4, on Charades-STA, our Mr. BLIP model achieves a new state-of-the-art improving over the previous SoTA model [24]. While [24] performs well on Charades-STA, it underperforms significantly on QVHighlights (e.g., about 17% drop on R1@0.7 compared to Mr. BLIP), highlighting the generality of our approach.
Furthermore, as shown in Table 4, Mr. BLIP outperforms the previous SoTA [58] in moment retrieval on ActivityNet Captions by 5.62% and 5.35% on R1@0.5 and R1@0.7, respectively.
Most previous works build upon features that were pre-extracted by convolutional neural networks [31], such as I3D [7] or Slow-Fast [19], often in combination with CLIP [49], with one recent state-of-the-art model, UnLoc [58], being the first to build solely on the image-text pretrained CLIP. Our work is most comparable to SeViLa [62], which leverages the same image-text pretrained MLLM, namely BLIP-2 [36], as its backbone. Our Mr. BLIP model achieves performance that is 20.14% and 23.29% higher than SeViLa on QVHighlights in Recall@1 at thresholds of IoU=0.5 and IoU=0.7. This demonstrates the effectiveness of our novel multimodal sequence design. We provide contextual and temporal information to our model and have it predict the relevant moments in a single, autoregressive pass-through. In contrast, SeViLa classifies each frame as relevant to the query or not, without any contextual and temporal information about the video.
To the best of our knowledge, ours is the first work casting the task of moment retrieval as an open-ended sequence-to-sequence problem and leveraging a generative MLLM that achieves the new state-of-the-art results on moment retrieval, which has been largely dominated by complex task-specific architectures.
4.4 Qualitative Results
Figure 3: Qualitative results. We showcase several example predictions from our Mr. BLIPRefer to caption model, including success and failure cases. Example 1 is from Charades-STA and examples 2-5 are from QVHighlights. Discussion in the main text.
In this section, we analyze qualitative results illustrated in Fig. 3.
In examples 1, 2, and 3, we observe the ability of Mr. BLIP to successfully localize one or multiple moments given a natural language query. Example 4 represents a case where our model predicts two correct moments alongside one that is false. Example 5 shows a failure case, in which our model predicts a moment where the woman is clearly visible in a shop but not trying on a hat. In the ground truth segment, the woman is trying on a hat but is not clearly visible since she is far away.
5 Conclusion
We introduce Mr. BLIP[Uncaptioned image], a multimodal large language model that performs moment retrieval by generating a variable number of relevant windows as a single sequence of text tokens given multimodal inputs. We show that Mr. BLIP does not require expensive video pretraining and greatly benefits from an input design that incorporates visual and temporal information as a simple interleaved token sequence of visual and text tokens. Mr. BLIP achieves state-of-the-art results on the most common moment retrieval benchmarks while leveraging a highly versatile design that allows for simple extension to further tasks. We hope this versatile sequence-to-sequence design of Mr. BLIP sparks further research in leveraging large-scale image-text pretrained MLLMs for computer-vision-based video understanding tasks.
A potentially interesting area to analyze is the impact and utility of the task prompt. It might be worth exploring training our model to handle multi-task problems and defining the output of Mr. BLIP via the task prompt.
5.0.1 Limitations and Impact.
Since we leverage a large-scale pretrained MLLM, BLIP-2 [36], we can not rule out the presence of any biases or harmful stereotypes (e.g., gender or racial biases) that BLIP-2 has learned. It is conceivable that our model will detect moments that fall into the category of harmful stereotypes. We recommend applying our model with caution when adopting it in practice. Moreover, our model could be subject to misuse w.r.t some forms of surveillance.
5.0.2 Acknowledgements.
As part of their affiliation with TU Darmstadt, the research of BM and MR was funded, in part, by a LOEWE-Spitzen-Professur (LOEWE/4a//519/ 05.00.002-(0010)/93) and an Alexander von Humboldt Professorship in Multimodal Reliable AI sponsored by Germany’s Federal Ministry for Education and Research.
Appendix 0.A Implementation Details
0.A.1 Video processing
We build on top of the Salesforce lavis repository and leverage their implementation of randomly and uniformly extracting frames from a video. During training, we randomly sample frames as a means of data augmentation to alter the frames seen by the model in each batch. At inference, we sample uniformly to provide an equal coverage of the entire video. For random sampling, we start by uniformly extracting n+1 timestamps, where the first and last timestamps are at $t=0$ and $t=video\_length$, respectively. Afterward, we randomly sample one frame between each pair of adjacent timestamps, yielding the final n randomly sampled frames. This process of random sampling can be considered adding noise to a uniform sampling process.
During training, we randomly crop and resize each frame to 224x224 pixels. We then normalize each frame by subtracting by a fixed mean and dividing by a specified standard deviation. At inference, we don’t apply random cropping and resizing. We only apply the normalization.
0.A.2 Hyperparameters
Table 5: Hyperparameters for our base models for Charades-STA [20], QVHighlights (QVH)[33], and ActivityNet (ANet) [29, 4].
Hyperparameter Charades-STA QVH/ ANet
Batch size 32 32
Epochs 50 50
Learning rate 3e-4 3e-4
Learning rate warmup Linear Linear
Learning rate warmup steps 10% 10%
Learning rate decay Cosine Cosine
Optimizer AdamW AdamW
Weight decay 0.05 0.05
# input frames 20 60
# beams 5 5
Table 5 shows a detailed list of hyperparameters. To ensure a fair comparison for all our models and ablations, we stick to hyperparameters presented in the table, except for the cases where we ablate certain parameters. The only hyperparameters that change across benchmarks are the number of steps for the learning rate warmup and the number of input frames. The number of warmup steps depends on the dataset size and corresponds to 10% of the total number of steps, i.e.
#$\#steps_{warmup}=0.1*\#steps/epoch*\#epochs$ (2)
Appendix 0.B Additional Results
0.B.1 Effect of number of trainable parameters
Table 6: Ablation on the number of trainable parameters.
LoRA Rank (# trainable parameters) R1@0.5 R1@0.7 mIoU
2 (6M) 66.03 43.03 56.59
4 (10M) 65.71 43.59 56.78
8 (19M) 67.28 46.70 57.46
16 (37M) 67.72 45.38 57.35
In the main paper, we state the use of parameter-efficient finetuning techniques and thus training only a small number of parameters. In practice, we leverage LoRA [22] with a rank set to 8 and apply it to all linear layers in the large language model, yielding about 19 million trainable parameters.
In Table 6, we show the effect of different LoRA ranks and their resulting number of trainable parameters. We observe that setting the rank to 16, i.e. training double the number of parameters does not yield any significant improvement (row 4). Moreover, training fewer parameters, i.e. 10 or 6 million, by setting the rank to 4 or 2, respectively, does degrade the performance of Mr. BLIP (rows 1 and 2).
Appendix 0.C Additional Qualitative Results
In this section, we provide further qualitative results and discuss shortcomings of our model that can be addressed in future research. For each example, we provide the discussion in the respective caption for the convenience of the reader. Examples 1 through 6 are from QVHighlights [33] and examples 6 through 10 are from Charades-STA [20].
As in the main paper, we illustrate the ground truth targets as dashed lines and the predicted windows as solid lines.
Figure 4: We observe the capability of our model to recognize and differentiate 3 separate moments that repeatedly occur with intermittent interruptions. Each predicted moment depicts the respective natural language query.
Figure 5: Mr. BLIP can recognize a distinct public figure, Donald Trump, although the training set includes only 6 queries containing Donald Trump.
Figure 6: Our model predicts a third window, that does not align with a ground truth moment, which, in fact, is accurate and depicts the query.
Figure 7: Mr. BLIP accurately predicts 3 out of 4 moments. It fails to predict the first moment, which lasts for only 2 seconds. This is likely because we sample a frame every 2.5 seconds and, by chance, might not sample one in the respective window.
Figure 8: The video contains two groups of relevant moments which contain a short sub-2-second cut that does not depict the query. With a resolution of 60 frames for a 150-second long video, corresponding to a frame being seen every 2.5 seconds, Mr. BLIP can not detect the cut and predicts two long moments that encompass the two short ones. Moreover, our model predicts a third window (red) that does not match a ground truth label. The predicted moment does in actuality depict soldiers, but those are not escorting other people.
Figure 9: Our model has trouble predicting very short moments in long videos and high-frequency jump cuts given the resolution of the frame sampling.
Figure 10: Given a very ambiguous query, our model’s prediction aligns with the ground truth.
Figure 11: Our model fails to recognize the action of the man closing the front door. We hypothesize this is because the door is very dark and poorly visible. Our model predicts the moment when the door is clearly visible and closed.
Figure 12: Our model fails to recognize the action of putting down the broom at the beginning of the video and predicts the moment when the man picks the broom back up and then holds it.
Figure 13: Mr. BLIP recognizes the moment when the person is sneezing. The ground truth is false.